### v2

# Characterization of NBI-driven shear Alfvén waves in the TJ-II stellarator using Mirnov probes and electrostatic potential fluctuation measurements

Abstract
We present the first experimental measurements of the toroidal mode number of shear Alfvén waves in the TJ-II stellarator. A series of experiments were carried out in three different magnetic configurations to investigate counter-NBI driven modes. Co- and counter- electron-cyclotron current drive was used to modify the rotational transform () profile leading to the destabilization of a varied set of Alfvén eigenmodes with different frequencies and mode numbers. To characterize the spatial structure of the modes we have used two Mirnov probe arrays, one dedicated to the measurement of the poloidal mode number and the other, a recently commissioned helical tri-axial array, dedicated to the measurement of the toroidal mode number. A heavy ion beam probe, operated in radial sweep mode, was employed to characterize the radial location of the modes. We show that the induced changes in , that are fundamental when it comes to validation studies, cannot be measured experimentally with motional Stark effect so, instead, the shielding current diffusion equation is solved in cylindrical geometry to estimate these changes. We calculate the incompressible shear Alfvén continuum for selected cases using STELLGAP and find reasonable consistency with observations. A database with the observed modes has been created, so that it can be used in future work for theory validation purposes.
bib/bib.bib
1 Introduction
The validation against experimental results of Alfvén Eigenmodes (AEs) physics models and of the numerical tools that implement these models is indispensable for the informed design of future reactors [fasoliChapter52007]. The experimental determination of the spatial structure of AEs through the identification of its poloidal (m) and toroidal (n) mode numbers and the measurements of the radial profiles of the electric potential is an integral part of this task.
The excitation of fast-particle induced Alfvén Eigenmodes has been routinely observed in TJ-II [cappaStabilityAnalysis2021, nagaokaMitigationNBIdriven2013, melnikovDetectionInvestigation2018]. NBI-driven modes measured in TJ-II plasmas appear as intense, narrow frequency bands in the spectrogram of magnetic, plasma potential or radiation fluctuations allowing, in principle, to have a clear diagnosis of the properties of the modes.
The main objective of the experiments described in this paper was to excite AEs in different scenarios and study their spatial structure, focusing on the mode number identification () by means of poloidal and helical arrangements of Mirnov coils [ascasibarMeasurementsMagnetic2022] on the one hand and, on the other hand, on the measurement of the radial profile of perturbed potential using Heavy Ion Beam Probes (HIBP). The measurement of the poloidal mode structure is routinely done since the poloidal array of Mirnov coils was commissioned and the determination of the poloidal number has been employed in previous studies [jimenez-gomezAlfvenEigenmodes2011, vanmilligenMHDMode2011]. Only very recently we have added the possibility to determine the toroidal mode number [ascasibarMeasurementsMagnetic2022] and the polarization of the wave detected at the coils locations. Beyond the interest related to model validation, the difficulty of this measurement, in particular in non-axisymmetric configurations, calls for a complete experimental study where we know that we are varying the spectrum of possible modes and can thus confirm that these variations are reflected in the helical array data and the toroidal mode numbers reconstructed from this data. Both measurements, magnetics and HIBP, along with the recent development of a synthetic diagnostic for Mirnov coil arrays [pons-villalongaExploringOperational2024] allow us to establish appropriate criteria to support the validity of the mode analysis results.
On-axis Electron Cyclotron Current Drive (ECCD) was chosen as actuator to modify the spectrum of AEs due to its minimal impact on the plasma density and temperature profiles compared to the effect it has on the plasma current, which isolates the rotational transform profile () as the driver of changes in the shear Alfvén wave (SAW) spectrum [cappaStabilityAnalysis2021]. To obtain a comprehensive database of different excited AEs, we conducted a parameter scan on magnetic configuration and EC beam injection angle. An essential piece of information when performing validation tasks is the rotational transform profile. Unfortunately, as we will show here, the changes induced by ECCD or Neutral Beam Current Drive (NBCD), which are important from the point of view of their impact on the Alfvén continuum, are not intense enough to be measurable by iota diagnostics such as Motional Stark Effect (MSE) [mccarthySpectrallyResolved2015]. Therefore we must rely on an estimation of the different external current sources and simulate the time evolution of rotational transform by calculating the radial diffusion of the shielding current.
This paper is organized as follows: section 2 contains a description of the experimental set-up and the relevant plasma parameters that were achieved in the experiments. Then, in section 3 we describe the mode analysis technique and the whole set of measured mode numbers. Next, in section 4, after a brief description of the HIBP measurements, we present several examples of the perturbed electrostatic potential profiles associated with the modes. In section 5, we solve the shielding current radial diffusion equation in cylindrical geometry taking the total experimental current as boundary condition and approximate profiles for the external current sources. This allows us to estimate the evolution of the rotational transform profile. A short discussion on the difficulty of resolving changes in using MSE measurements is also included in this section. Although the main purpose of the paper is the experimental characterization of instabilities in different scenarios, in section 6 we will take a couple of cases and compare the observations with the Alvén continuum calculated with the STELLGAP code [spongShearAlfven2003]. Finally, we come to our conclusions and lay the future groundwork for the implementation of the validation studies.
2 Experimental arrangement and results
TJ-II [ascasibarOverviewTJII2001] is a four period medium-sized stellarator of the heliac type with major radius m, minor average radius m, magnetic field on axis T and plasma volume m3. NBI sources produce co and counter tangential beams at 32 kV. Since counter injection with NBI2 produces a richer spectrum of modes and NBI1 driven AEs have been already the subject of research in a previous paper [cappaStabilityAnalysis2021], the present study focuses only on the analysis of the modes produced by the NBI2 injector. Figure 1 shows the launching configuration around the device of both neutral beams and the two electron cyclotron heating beams (second harmonic X-mode).
To investigate the spatial structure of the AEs destabilized in different magnetic configurations and with variable on-axis ECCD settings, the following experimental set-up was used. The heating sequence is shown in figure 2. The plasma is started using ECRH, with both beams switched-on simultaneously. One of them is set to inject perpendicularly to the magnetic axis (ECRH2) and the other (ECRH1) is launched with different incidence angles that are varied on a shot to shot basis. After ms, NBI2 is switched on, creating a fast-particle population that heats the plasma when slowing down and at the same time induces current (NBCD) and drives the AEs. The ECRH2 beam with perpendicular incidence is turned off shortly after the injection of NBI2, while the ECRH1 beam remains on to both drive ECCD and help to maintain a constant plasma density during the rest of the shot.
We have performed the same scan using three distinct magnetic configurations with different values of rotational transform on-axis and very similar low shear. For our present purposes, although the central rotational transform can be varied between and , we will refer to them as low (), medium (or standard, ) and high iota () configurations. The vacuum profiles are represented in figure 3, along with a poloidal cross section of the magnetic surfaces of each configuration at .
Exploring configurations with large or lower values of is generally more demanding in terms of plasma performance and density control. In this case, except for the high configuration that needed kW to improve density control, the rest of shots were achieved with kW, kW and kW. As was mentioned in the introduction, the advantage of using on-axis ECCD to modify the spectrum of shear Alfvén waves is that the small changes induced in the plasma density and temperature profiles when steering the beam along the axis have a smaller incidence on the SAW spectrum than the changes produced by variations in the toroidal plasma current. Five positions of the ECRH1 beam were used to drive different amounts of toroidal plasma current. They are labeled according to the on-axis wave parallel refraction index in the plasma disregarding refraction effects, , where is the angle between the wave vector of the injected wave and the direction of the magnetic field along the axis.
-0.2 | -0.1 | 0 | 0.1 | 0.2 | |
[∘] | 101.5 | 95.7 | 90 | 84.3 | 78.5 |
Negative values of correspond to co-ECCD, generating positive current (increasing ), while positive ones stand for counter-ECCD, that generates negative current (decreasing ). Table 1 shows the ECRH1 injection angle in degrees for each value of .
Figure 4 compares the time traces of the line densities when this type of scan is performed. As expected, line densities practically overlap while plasma current variations up to 3 kA are measured at the end of the shot.
This behaviour is illustrated in figure 5. The plasma current exhibits low frequency oscillations originated by a small ripple in the current that flows through the central and helical main coils of the device. As we will see later in section 5, following a procedure detailed in [mulasValidatingNeutralbeam2023], this inductive contribution can be extracted from the Rogowski measurement and a “clean” current may be obtained. Counter injection with NBI2 produces negative NBCD and therefore maximum negative current is achieved when counter-ECCD is applied.
Figure 6 shows the plasma density and temperature profiles in the medium iota configuration, reconstructed by integrated Bayesian analysis of Thomson scattering data (taken at ms) and profile reflectometry data [vanmilligenIntegratedData2011]. Note that they also overlap within the error bars confirming that the changes brought about by the ECCD barely impact the thermal plasma profiles.
A good density control is exhibited during the time interval of interest (between and ms) for the low and medium configurations. This simplifies the mode number analysis, allows for reproducible measurements of mode potential radial structure and is fundamental when carrying out validation studies, that benefit from having a steady state fast ion distribution and a steady state NBI2-driven current source. In the case of the high configuration, density control was harder to achieve due to degradation of wall condition, requiring an increase in ECRH1 power during NBI2 phase (up to 200 kW) and also a decrease in pulse length.
3 Magnetic fluctuations and mode number measurements
There are two arrays of Mirnov coils, that digitize the magnetic fluctuations at a rate of 1 MS/s. One of them, hereinafter referred to as poloidal array, consists of 25 single-axis coils distributed in a vertical plane at constant toroidal angle, and provide information on the poloidal mode number by measuring magnetic fluctuations in the (approximate) poloidal direction [jimenez-gomezAlfvenEigenmodes2011, jimenez-gomezAnalysisMagnetohydrodynamic2007]. The other array, hereinafter referred to as helical array, is distributed toroidally along a full period of the device, and is comprised of two up-down symmetrical sub-arrays of 32 tri-axial sensors each [ascasibarMeasurementsMagnetic2022] that follow the helical path of the plasma. These arrays are used to identify both toroidal () and poloidal () numbers of the excited modes. Additionally, there are also two vertical arrays located at different toroidal angles not used for mode periodicity measurements. The mode analysis is carried out using the Lomb periodogram [lombLeastsquaresFrequency1976, zegenhagenAnalysisAlfven2006], that can handle non-equispaced data points. An in-depth discussion on the mode analysis can be found in [pons-villalongaExploringOperational2024], so here it will just be outlined. For the mode analysis, each coil has to be mapped to a pair of magnetic angles, that in our case will be the Boozer angles of the closest point to the coil in the Last Closed Flux Surface (LCFS). These angles are found using the vacuum equilibrium of each configuration, as tests with VMEC equilibria calculated with different plasma current profiles show that their variation is negligible in comparison with the error induced by the uncertainty in the positioning of the coils. Figure 7 shows the distribution of the poloidal and helical arrays around the plasma and the Boozer angles for the standard configuration.
3.1 Observed eigenmodes
The spectrum of NBI2 driven modes observed in each of the three magnetic configurations, for all the values of explored, is shown in figure 8. The fact that the plasma density is not exactly constant throughout the shot allows us to identify the Alfvénic nature of the excited modes by their slow evolution in frequency, an effect which is largely amplified when the ECRH1 heating shutdown at ms produces a sudden increase in plasma density and a concomitant reduction in mode frequency ( as expected for AEs). This figure also illustrates how the spectrum of modes is modified as the ECRH1 beam is steered along the magnetic axis. In the medium case, the shear Alfvén spectrum clearly evolves from a situation with three main branches well separated in frequency to a situation in which only a single intense mode is detected. Similar changes are observed in the high case while almost no modes are observed for the low configuration during the ECRH1+NBI2 phase. This set of shots shows a wide array of behaviours of the magnetic fluctuations. Among the most remarkable we find:
Modes that emerge and vanish, for example in shots 53491 and 53487, where transient modes can be observed above the kHz mode that is present during the entire NBI2 phase. This could be caused by the strong low frequency inductive oscillations in current mentioned above and the associated changes in iota or by non-linear energy transfer between the modes, both effects being outside the scope of the paper.
Transitions between chirping and stable phases, and viceversa, seen in shots 53806 (and 53807, not shown here). In both cases, the toroidal current reached at the end of the shot is maximum and negative while it also exhibits the fastest time evolution, both observations consistent with the transition from chirping to steady frequency mode [melnikovTransitionChirping2016].
Coupling of modes well separated in frequency, as happens in shot 53487, where the mode at kHz becomes more intense at ms, while at the same time the mode at kHz weakens.
An example of chirping mode, analyzed via Short-Time Fourier Transforms (STFT) and the DMUSIC method [kleiberModernMethods2021] is shown in figure 9. The mode displays rapid chirping behaviour, with a repetition frequency of about 1 kHz. These chirping modes are more difficult to analyze in terms of mode numbers, as the rapid sweep in frequency deteriorates the precision of the analysis technique discussed below. Aditionally, figure 10 shows again a comparison of STFT and DMUSIC spectrograms calculated for shot 53490. A closer look tells us that what looks like a single intense mode is actually formed by three distinct frequency components separated by about 4 kHz from each other, and that remain remarkably stable during the considered time interval. The DMUSIC method is very useful in these cases, as it helps in the precise determination of the mode frequency, that is needed for the periodogram. Figure 10 also shows the bandpass filtered signal (between 250 and 310 kHz), that shows large amplitude oscillations interleaved with periods
of lower amplitude as corresponds to a combination of three near-frequency modes of similar amplitude. Despite the clear frequency structure, the analysis cannot resolve different mode numbers for each of the observed frequencies. Even applying very narrow bandpass filter intended to isolate each frequency contribution the mode number result remains the same (see section 3.2).
3.2 Mode number identification
The Lomb periodogram of a set of signals , where is a time index and a coil index, is given by
where is the phase term, is a normalization factor, and is a phase shift term, given by
that makes the periodogram invariable to time and angular shifts.
The Lomb periodogram requires as input the frequency of the mode, that is obtained from the spectrogram, and the integer pair of mode numbers (), noted to as hereinafter, over which a scan is carried out to obtain a two-dimensional distribution of intensities. Given the direction chosen for the coil distribution arrangement and the TJ-II boozer coordinate system, the sign criterion (+) in the phase term is such that a mode number with positive and positive propagates toroidally along the direction of the magnetic field and poloidally along the direction of the ion diamagnetic drift.
After determining the real orientation, and correct phase sign of each set of coils by means of dedicated calibration experiments described in [ascasibarMeasurementsMagnetic2022], the measured can be projected over any desired polarization vector basis (only in the case of the helical array). In this work we will use the projection of along the normalized tangent basis vector , also in Boozer coordinates, as it proved to be the best-performing basis for the mode number analysis [pons-villalongaExploringOperational2024]. A suitable collection of time and frequency intervals are selected for each mode, and the signals, calibrated and expressed in the chosen polarization basis, are then filtered and normalized to minimize distance effects on the periodogram result.
The signals from the poloidal and helical arrays cannot be included in the same periodogram, as a difference in coil orientation induces a phase shift in the signal that compromises the accuracy of the analysis. To work around that, periodograms for the poloidal and helical arrays are computed independently and then multiplied. Furthermore, this method also helps work around the main limitation of the helical array, that is, that the coils follow a straight path in Boozer coordinates so, in an aliasing-like phenomenon, a strip of pairs of mode numbers with slope , that contains the true pair of mode numbers, appears in the periodogram. The product with the periodogram of the poloidal array, that suffers to a much lesser extent from the same issue, enables in principle the determination of the true pair.
Figure 11 shows an example of this process applied to a low frequency mode observed in one of the medium configuration shots, with the periodograms for the poloidal array (left), helical array (center) and their product (right). The identified pair of mode numbers is . As we have already mentioned, a fundamental drawback due to the non-axisymmetric configuration and the spatial distribution of the coils is that there are still modes with different pairs of characteristic mode numbers that appear with almost the same probability in the analysis and we must use arguments based on the expected values of the rotational transform to discard some combinations of modes versus others. Low frequency MHD modes have shown to be easier to identify than AEs with higher frequency even though they exhibit larger bandwith, as is the case of the modes analyzed in this work. The quality of the mode analysis deteriorates with rapid changes in frequency and mode intensity. Furthermore, systematic errors during data acquisition and reduced coil availability cause a broadening of the mode number strips and sometimes introduce false positives. High frequency broadband modes are also difficult to analyze using this method, and require more sophisticated techniques such as the Stochastic System Identification (SSI) [kleiberModernMethods2021]. Especially deleterious for the analysis are spurious phase shifts [horvathReducingSystematic2015], that are also difficult to characterize experimentally for the coil arrays of TJ-II.
Typically, the difference in amplitude between the most intense pair and the next is small, as can be seen in figures 12 and 13, making a single analysis per mode insufficient for a confident mode identification. To circumvent this issue, many periodograms have been conducted per mode (see figure 14 for an example), selecting suitable time-frequency intervals over a spectrogram. This process has been conducted manually in this work but can be automatized using mode tracking algorithms [vazmendesBroadbandAlfvenic2023, heinrichCharacteristicsAlfvenic2024, bustosAutomaticIdentification2021].
Then, for each mode, the most intense pairs in each PAHA periodogram are selected, and the dominant mode number is identified as the most frequently occurring pair. With this approach, we aim to reduce the uncertainties in the mode identification and improve the reliability of the results, while at the same time proceeding systematically.
Figure 15 shows the aggregate results of the analysis, for each configuration and . In some cases, the variability in the most intense mode number is too large to identify the dominant mode number, so we apply a cutoff based on the relative frequency of ocurrence of the most frequent mode number (), in the following way:
High confidence
Low confidence, but worthy of consideration after manual review
Mode could not be identified
In other cases, the most frequent mode number pair lacks clear physical significance, so in those cases the second most common pair is selected, provided it appears as the most intense mode a comparable number of times. That this can happen should not be surprising, especially after inspecting figures 12 and 13. These instances are marked in figure 15 by coloring the mode numbers in red. Finally, some modes are located too close to the core to be identified [pons-villalongaExploringOperational2024]. In these cases, in which HIBP measurements can be used to distinguish true low-, low- modes from core-localized ones (more in the next section), the mode numbers are colored blue. A contains tables with the most frequent dominant mode numbers for each scenario.
4 HIBP Measurements
Two HIBP systems (see figure 16) are installed in TJ-II [bondarenkoInstallationAdvanced2001, melnikovHeavyIon2017], each beam probe located in a different toroidal position.
They provide data on the plasma potential, plasma density and poloidal magnetic field in the beam sample volume, that is usually swept along the plasma radius ( ms) so as to provide profiles of these quantities. Of particular interest for this work are the plasma potential perturbation profiles, that can be used to estimate the radial positions of the measured AEs. In this case HIBP1 was used in radial scan mode while the other probe (HIBP2) was only operated in some of the shots in fixed point modality, measuring plasma potential at .
Figure 17 shows an example of the measured quantities for the case in the medium configuration (shot #53492). The spectrogram of the magnetic fluctuations corresponding to a shot with the same conditions (#53491) is shown in figure 8. The more or less stationary plasma, which translates into a constant mode frequency, allows us to average over repetitive scans of the heavy ion beam and retrieve a radial profile of the mode. This result, that can already be guessed by inspecting the spectrogram of plasma potential fluctuations presented in figure 17 is shown in figure 18.
In this case, the spectrogram of magnetic fluctuations shows three distinct modes around 180, 215 and 310 kHz. The one with lower frequency is barely seen by HIBP1 in either density or potential fluctuations. The radial profile measurements reveal a more core localized 215 kHz mode with an approximately symmetrical profile while the 310 kHz mode is shifted towards the medium region of the plasma and exhibits of somewhat asymmetrical profile. The corresponding mode numbers extracted from the Mirnov coils data were given in figure 15 and appear also in the figures. The bottom panel in figure 18 shows the dependence of the secondary beam intensity (Cs++) on the radial position. As shown in reference [khabanovDensityProfile2019], the density profile, roughly proportional to the beam intensity profile can be reconstructed from this measurement. Note that, consistently with the density profiles shown in figure 6, the intensity gradient region falls between and . This is important when considering the excitation of low frequency MHD modes and its relation to the rotational transform profile that will be explored in section 5.
As mentioned above, for some modes a radial asymmetry in the potential profile is clearly observed. This is consistent with the fact that these are probably gap modes [kolesnichenkoAlfvenEigenmodes2002]. These modes consist of coupled modes with different ’s and ’s that evolve together. The difference in poloidal mode numbers produces the asymmetry in the potential profiles. On the other hand, as it was demonstrated in reference [pons-villalongaExploringOperational2024] using a synthetic Mirnov coils diagnostic, in the case of coupled modes, the magnetic coils cannot resolve different ’s when the difference between them is or . In these case, only the dominant mode numbers result from the analysis. The radial profiles of the mode potential have been measured for several values of and additional results for and for the other magnetic configurations are shown in figures 19, and 20, 21 respectively.
5 Time evolution of rotational transform
To understand the changes that occur in the Alfvén wave spectrum when we modify the direction of the microwaves beam it is necessary to have an estimate of the time evolution of the rotational transform. When this evolution is due to a current source that is established from almost the beginning or when we observe that the value of the toroidal current measured by the Rogowski coil stabilizes before the end of the shot we can assume that, in the final part of the shot, the plasma shielding current is zero and the current profile inside the plasma is that of the current source that we force in the plasma, either using ECCD or NBCD, or is the one produced by the bootstrap current inherent to the configuration and the plasma profiles. With this approximation we can have an estimate of the rotational transform at the end of the shot [cappaStabilityAnalysis2021]. In this case this cannot be done because we have a source of ECCD current during the whole shot duration, and a second source due to NBI2 which is established a few milliseconds after the neutral beam injection. Furthermore we observe in most of the cases that the current is not stabilized at the end of the shot.
5.1 Radial diffusion of toroidal current
The time evolution of the toroidal current and consequently of can be estimated using a simple model of the radial diffusion of the different current sources in a cylindrical plasma. We start from Maxwell’s equations where we neglect the displacement current term,
and we assume that the toroidal current is related to the electric field through Ohm’s law,
In this case the toroidal current source is a combination of the current induced by ECCD, NBCD and bootstrap. In these low density plasmas the bootstrap current is smaller than the currents due to the heating systems, it is not centered at the core, its impact on is therefore smaller, and we will not take it into account in the calculation. Also, we assume that the ECCD and NBCD current sources are not time-dependent. This assumption is reasonable since we have an approximately constant density throughout the shot. In addition, we also assume a constant conductivity over time, but which may not be the same depending on whether we are in the ECRH or ECRH+NBI phase due to changes in plasma temperature. In fact, as we see shortly, the complete time evolution must be calculated in two phases whenever an ECCD contribution ( is present.
Combining equations 3 and 4, assuming only radial dependencies of all quantities and assuming also vector components only along the direction parallel to the static field , we arrive at the following equation for the time evolution of the shielding current ,
where is the plasma resistivity and is the integral of the shielding current density due to the plasma response,
Equation 6 is equivalent to the one given in [zaniniECCDinducedSawtooth2020] taking into account that the relation between the current and iota in the cylindrical case is given by
For a cylindrical plasma with minor radius , central resistivity and time constant , we can introduce dimensionless variables , and and write equation 6 as
where we have defined . We solve this equation numerically with a standard MATLAB ® solver. Once is known we may write the total rotational transform as
5.2 Initial and boundary conditions
According to expressions 5 and 7, the shielding current is related to the total plasma current as follows
and therefore the initial and boundary conditions for are given by
As for the initial conditions, we stated above that the main current sources come from ECCD and NBCD, that is . For both the ECCD current at different injection angles and the NBCD current produced by the counter injector we will use a combination of estimates presented in previous works together with the plasma toroidal current measured in the experiments discussed here.
The NBCD profile has been taken from the ASCOT5 slowing-down simulations presented in [mulasASCOT5Simulations2022]. The total NBCD current obtained in the medium configuration was validated against experimental results in [mulasValidatingNeutralbeam2023] and a good agreement was observed for neutral beam counter-injection at line densities and density profiles similar to the ones shown in figures 4 and 6. For our present purposes and having in mind that different configurations are used and that the plasma densities slightly differ between the three of them, the integrated current of the NBCD source will be determined from the plasma toroidal current measured in the cases and we will take the current density profile calculated in reference [mulasValidatingNeutralbeam2023]. We mentioned in section 2 that the low frequency oscillations originated by inductive coupling with the main conductors can be extracted from the Rogowski coil measurement. Note that these oscillations not only occur with respect to a certain average value, but may also be shifted upwards or downwards depending on the specific behavior of current ripple in the central and helical coils of the device. Applying the procedure described in the appendix of [mulasValidatingNeutralbeam2023] to the current measured in the NBI2 phase we arrive at the result shown in figure 22.
The values of the integrated NBCD current source (no ECCD in this case) in the limit of zero shielding current and the LR time for each case are given in the figure. The current measured in the medium and high iota configurations (actually the limit current inferred from the measurements) agrees well with the results presented in [mulasValidatingNeutralbeam2023]. However, the NBCD measured in the low iota configuration appears to be lower. A possible reason for this is the slightly lower plasma volume and size of this configuration (see figure 3) and also the fact that plasma line densities in this case () were a bit lower than the line densities achieved in the other two (). The beam shine-trough power and subsequent slowing-down ASCOT5 simulations that would clarify these differences are out of the scope of this work.
The ECCD profile is estimated from experimental observations of ECCD total induced current in oblique launching experiments [fernandezECCDExperiments2008] and fast power modulation experiments [eguiliorHeatWave2003] for perpendicular injection combined with the results of ray-tracing and beam tracing simulations [martinez-fernandezHighPower2020]. We use these simulations to determine how the power deposition and ECCD profiles widens as increases. This result is corrected with the measured power deposition profile at by imposing that its integral is the total measured current after several LR times. With these assumptions we obtain the current densities that are shown in figure 23.
Regarding the boundary condition at , we see that the quantity is nothing else than the experimental toroidal current measured by the Rogowski coil that is shown in figure 5. For stationary current sources, the evolution of the toroidal current can be expressed as
and the boundary condition takes the final form
Since all of the above is only valid for stationary current sources, the complete time evolution when must be calculated in two phases; first with an ECCD current source starting from zero plasma current and then with a combined ECCD+NBCD current source starting from the final plasma current profile given by the first simulation. Note that since , the time constant of the plasma column () is actually determined experimentally by fitting the measured time evolution of the plasma current to its theoretical evolution given by expression 11. In this case, following the definition of given above, this is equivalent to determine an experimental central conductivity for the cylindrical plasma. We do not intend to compare this result with the actual plasma conductivity that may obtained from neoclassical theory since neither the inductances of both infinite cylindrical and real plasma columns, nor the expressions for , are the same.
5.3 Iota evolution results without ECCD
The result of applying this model to the cases with no ECCD current () are shown in figure 24.
The calculated evolution shows that for these cases in which the toroidal current is still evolving (see figure 22), the profile of at the end of the shot ( ms) is midway between the vacuum value and the equilibrium value. It is enlightening to compare the profiles shown in the figures with the mode numbers of the low frequency MHD instabilites ( kHz) shown in figure 15. Restricting ourselves to the shots with for the moment, the mode number analysis shows a instability for the low configuration, a for the medium case and two instabilites with mode numbers and for the high case. This is consistent with the calculated profiles considering that this type of instability is triggered in the density gradient region () in the presence of a low order rational value; , and , for the low, medium and high respectively. Note that following the numerical study of Mirnov coils performance presented in [pons-villalongaExploringOperational2024] the calculation of the mode numbers is subject to an indeterminacy of or greater depending on the radial position of the instability and thus the pairs and delivered by the analysis may very well correspond to and having in mind that actually and are well outside the reasonable values when negative NBCD current is at play.
5.4 Iota evolution results with ECCD
Let us now look at how this low frequency activity changes as a function of the ECCD current amplitude. Take for instance the mode observed in the low case. For even more negative current, that is, adding counter-ECCD with to the already negative NBCD, the mode is always present with more or less amplitude (that probably depends of the position of the rational in respect to the density gradient).
For maximum positive ECCD (), disappears since positive current elevates the profile. The most interesting case, for which the calculation of the time evolution of iota may give us a clue, is the one for . In this case, co-ECCD pushes-up the profile but at soon as NBCD comes into play, its contribution makes the rational surface appear again. This is also the case for which the perturbation shows the largest amplitude in the magnetic signal. The two-step model cannot capture exactly the time sequence of the experiment but clearly illustrates the phenomenon. This is shown in 25, where both the final iota profile due to ECCD current only and the final profile due to the combination of both sources is presented. The first one does not contain the low order rational while the second one intersects at and .
With the current parameters of this cylindrical model we cannot reproduce quantitatively the experimental behaviour but the main trends are well illustrated. Note that the combination of both negative and positive current sources, one broader than the other, produce a minimum in the profile. Same reasoning can be applied to the low frequency instabilities observed for the other two magnetic configurations. Figure 26 shows a case of maximum negative current () for the medium configuration. The 8/5 rational surface is located around . Even for this maximum negative current the 3/2 rational is in the inner plasma and therefore no low frequency mode appears associated with it.
5.5 MSE measurements
At the time of carrying out these experiments, the MSE diagnostic was not available and the measurements presented here were performed a posteriori in the medium configuration only, using a very similar plasma density.
Plasma temperature was higher since, due to poor wall conditioning, the ECRH2 beam had to be maintained throughout the shot to help control the plasma density. MSE measures the orientation of the magnetic field vector perpendicular to the injection direction of the diagnostic neutral beam [mccarthySpectrallyResolved2015]. This information is usually given in terms of the angle , where are the positions along the beam where the lines of sight of the spectrometer intersect the beam. The basic set-up and the values of for the medium iota configuration in vaccum are shown in figure 27.
Using the VMEC code to calculate the plasma equilibrium and its profile for a given profile of toroidal current, and determining the components of the field perpendicular to the beam, we may estimate the quantity measured by MSE and establish a correspondance with the profile. Figure 28 shows the results of the MSE measurements for different launching angles of the ECRH1 beam. These were made at the end of the NBI2+ECRH(1&2) phase where both NBCD and ECCD are acting to modify the rotational transform. Although MSE can clearly distinguish between rotational transforms of different configurations [mccarthySpectrallyResolved2015], it cannot resolve changes in due only to different plasma currents.
Actually, this can be confirmed calculating the expected for different VMEC equilibriums modified by toroidal current. Figure 29 compares the profiles of and without plasma current and for core localized ECCD current densities with +2 kA and -2 kA. It is shown that even though the changes in current are relevant from the point of view of Alfvén modes excitation, they are not large enough to be detected by MSE, whose error bars greatly exceed the variations expected from changes in plasma current.
6 Shear Alfvén continuum with STELLGAP
Similarly to what happens in tokamaks, where modes of the same and different couple within the TAE gap, the existence of helicity gaps (HAE) in stellarators allows for complex radially extended and weakly damped structures characterized by modes with different values of and . The complexity of the periodogram results presented in section 3 is a clear evidence of this. STELLGAP [spongShearAlfven2003] solves the Alfvén continuum equation for 3D stellarator equilibria taking into account the interactions between multiple toroidal modes. The low incompressible limit is considered here and coupling to sound waves is not included. We have only analyzed a couple of cases, the ones belonging to the medium configuration that exhibit the most diverse activity, i.e. and (see figure 15). The corresponding rotational transform profiles calculated at ms are shown in figure 30.
Gaps in the continuum, that favors the existence of radially extended discrete modes, are classified according to the mode numbers of the coupled structures, and this, among other things, depends on the number of periods of the device. In a stellarator, a necessary condition for mode coupling is that the toroidal mode numbers and of two different modes differ by a multiple of the number of periods (). Modes that verify this condition belong to the same mode family, which are labeled according to the lower mode number of its constituents. For instance, modes with and couple in the TAE gap and modes with and couple in the HAEμν gap. TJ-II stellarator has four periods and this results in three distinct mode families. References [schwabIdealMagnetohydrodynamics1993, spongShearAlfven2003, kolesnichenkoAlfvenEigenmodes2002] provide a more in-depth treatment of the subject.
Figure 31 shows the result of the STELLGAP calculation for . Each plot corresponds to one mode family. The sign criterion for the phase term in STELLGAP () is opposite to the one used in the Lomb periodogram analysis and therefore must be replaced by when comparing the results of the code to the measurements. Two helical gaps (HAE21 and HAE31) in the spectrum appear for mode families and . TAE gap appears higher in frequency and is not shown here. According to the result shown in figure 15, the 9/4 mode, observed around 175 GHz, would correspond to the {, } pair in the HAE31 gap. Previous studies with a synthetic Mirnov diagnostic [pons-villalongaExploringOperational2024] show that, besides uncertainties in poloidal mode number determination up to , Mirnov coils arrays in TJ-II cannot distinguish different constituents in helically coupled structures and therefore we cannot rule out that such combination of modes may show as an when measured. Other modes observed in the case, identified with less degree of confidence as and could have its correspondance in the family modes. In this case, the minimum in frequency of the mode favors the presence of a global Alfvén eigenmode (GAE) while the mode would have to be an energetic particle mode (EPM) since it would intersect the continuum. None of this can be assured with certainty without calculating the profiles of the modes, using a code such as AE3D, and their destabilization rates by means of linear stability simulations using gyro-fluid or gyrokinetic codes such as FAR3d or EUTERPE respectively.
In the same way we can take now the rotational transform profile of the case and calculate the corresponding continuum.
The minimum in the rotational transform profile around , produced by the combination of positive ECCD and negative NBCD has a strong impact of the shear Alfvén spectrum as we could already guess from the experimental observation that shows up to five different modes in the magnetic fluctuations spectrum (see figure 14). Modes with low mode numbers , , determined with a high degree of reproducibility in the periodogram analysis, and very close in frequency are consistent with the continuum of the family shown in figure 32. There, maximums and minimums in the continuum, also close in frequency, and prone to GAE destabilization appear around with mode numbers and . Moreover, a high frequency mode is observed around 310 kHz, where the calculation for predicts also high mode numbers AEs in the HAE21 and HAE31 gaps that merge in the plasma periphery. HIBP measurements for these modes are shown in figure 19.
7 Conclusions
We have described a series of experiments performed in the TJ-II stellarator. ECCD was used to modify the profile in order to destabilize different NBI-driven Alfvén eigenmodes and to have a wide range of cases available for further validation studies. The observed modes have been diagnosed with two arrays of Mirnov coils, used to determine the toroidal and poloidal mode numbers, and one heavy ion beam probe operated in radial sweep mode to measure the radial profile of the perturbation. This is the first time that the toroidal mode number of Alvén waves in TJ-II has been measured. Significantly, radial asymmetries were observed in the potential profiles of some modes, which points to their character as gap modes. There remain uncertainties in the mode number measurements, some due to the inherent limitations of the diagnostic and others that could be mitigated with an improved analysis technique, such as the SSI method [kleiberModernMethods2021], or Dynamic Mode Decomposition (DMD) [schmidDynamicMode2010].
For the purpose of theory validation, the determination of the radial profile of the rotational transform is essential. It is shown that changes in iota that can lead to significant variations in the Alfvén continuum are not measurable with MSE and, instead, a cylindrical model for the evolution of the shielding current in response to the NBI- and EC-induced currents has been used to estimate it. The profiles obtained with this model have been used as inputs for STELLGAP, and the results show a reasonable consistency between the calculated shear Alfvén continua and the experimental observations. Only the cases with the most intense activity have been analyzed with STELLGAP, and other interesting cases, such as the one exhibiting a structure with three very close frequencies and a clearly asymmetric radial profile, all compatible with gap mode behavior, have been left for a more theoretical treatment.
The database of AEs obtained in different scenarios, characterized by their mode numbers and radial profiles of electrostatic potential can be used for theory validation purposes using gyro-fluid codes such as FAR3d [varelaStabilityOptimization2024], already applied to co-NBI driven modes in TJ-II [cappaStabilityAnalysis2021] or gyrokinetic codes such as EUTERPE [kleiberEUTERPEGlobal2024], that was applied to study AEs fast-ion drive in W7-X [slabyInvestigationMode2020]. Thermal plasma profiles and estimates of neutral density profiles are available and fast ion slowing down simulations with ASCOT5 in the three magnetic configurations can be performed to obtain the corresponding fast ion density profiles. In addition to mode potential, HIBP measurements can also provide the density fluctuations profile associated to the Alfvén mode [melnikov2DDistributions2022], which are very relevant when comparing with the theoretical predictions.
A more accurate estimation of the evolution of the rotational transform profile could be considered in the event that the theoretical predictions did not match the observations. It can be obtained by solving the current diffusion equation in the real TJ-II geometry [strandMagneticFlux2001, schmittModelingMeasurement2010]. Furthermore, this model assumes that both the ECCD and NBCD currents are established instantaneously at the time at which the ECRH or NBI beams are injected. This is an approximation, particularly in the case of NBI in which the current source needs a time of the order of the slowing-down time of the fast ions to fully form. The self-consistent calculation addressing the impact of NBCD on would involve the use of a transport code such as ASTRA coupled to ASCOT5.
8 Acknowledgements
The authors want to thank the TJ-II Team for their support. This work was partially funded by the Spanish Ministry of Science and Innovation under several contracts, Grant Nos. ENE2013-48109-P, FIS2017-88892-P, FIS2017-85252-R, PID2021-125607NB-I00, and by the European Regional Development Fund (ERDF) “A way of making Europe”. This work has been carried out within the framework of the EUROfusion Consortium, funded by the European Union via the Euratom Research and Training Programme (Grant Agreement No 101052200 — EUROfusion). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the European Commission can be held responsible for them.
Appendix A Mode number results
This appendix contains tables that summarize the mode number analysis. In them, only the modes that appear as most intense more than 15% of the time are shown, along with the total number of analyzed time intervals. If no mode appears the minimum amount of times, then no mode numbers are shown.
f [kHz] | () | Counts | |
-0.2 | 260.0 | — | — |
Total: 22 | |||
-0.1 | 17.4 | (-3, -2) | 14 |
(1, -1) | 5 | ||
Total: 21 | |||
-0.1 | 368.2 | (13, -15) | 8 |
Total: 45 | |||
0.0 | 13.0 | (-3, -2) | 12 |
Total: 12 | |||
0.0 | 273.2 | (-14, -3) | 29 |
Total: 33 | |||
0.1 | 12.7 | (-3, -2) | 10 |
Total: 14 | |||
0.1 | 251.9 | (6, 2) | 12 |
(13, 5) | 6 | ||
Total: 30 | |||
0.2 | 11.0 | (-3, -2) | 12 |
Total: 14 |
f [kHz] | () | Counts | |
-0.2 | 47.1 | (-9, -5) | 13 |
(-8, -5) | 3 | ||
Total: 18 | |||
-0.2 | 219.8 | (-14, -6) | 10 |
Total: 32 | |||
-0.2 | 153.6 | (-9, 1) | 5 |
(11, 4) | 5 | ||
Total: 30 | |||
-0.2 | 301.8 | (-1, 3) | 8 |
(13, -14) | 7 | ||
Total: 32 | |||
-0.1 | 17.3 | (-9, -5) | 11 |
Total: 12 | |||
-0.1 | 187.2 | (6, 3) | 13 |
Total: 15 | |||
-0.1 | 221.0 | (-1, 0) | 5 |
(0, 0) | 4 | ||
(-2, 0) | 2 | ||
Total: 13 | |||
-0.1 | 206.6 | (-3, -2) | 4 |
(1, -1) | 4 | ||
Total: 8 | |||
-0.1 | 165.0 | (4, 3) | 7 |
Total: 13 | |||
-0.1 | 301.6 | (13, -14) | 5 |
Total: 9 |
f [kHz] | () | Counts | |
0.0 | 20.3 | (-9, -5) | 28 |
Total: 31 | |||
0.0 | 261.1 | (5, 2) | 11 |
Total: 31 | |||
0.0 | 177.8 | (9, 4) | 22 |
(-3, 1) | 15 | ||
Total: 38 | |||
0.0 | 193.7 | (1, -1) | 11 |
(-3, 0) | 5 | ||
Total: 31 | |||
0.0 | 294.4 | (-12, -1) | 10 |
(6, 4) | 8 | ||
Total: 30 | |||
0.1 | 21.1 | (-9, -5) | 10 |
(-8, -5) | 4 | ||
Total: 21 | |||
0.1 | 275.3 | (12, 5) | 9 |
(0, 5) | 5 | ||
(-7, -1) | 5 | ||
Total: 32 | |||
0.2 | 21.6 | (-9, -5) | 8 |
(-8, -5) | 5 | ||
Total: 14 | |||
0.2 | 270.8 | (12, 5) | 10 |
(13, 5) | 7 | ||
Total: 19 |
f [kHz] | () | Counts | |
-0.2 | 38.5 | (14, 0) | 4 |
Total: 21 | |||
-0.2 | 192.8 | (-10, -1) | 7 |
Total: 41 | |||
-0.2 | 271.1 | — | — |
Total: 24 | |||
-0.1 | 55.6 | (-9, -5) | 14 |
(-8, -5) | 3 | ||
Total: 19 | |||
-0.1 | 191.1 | (-15, -5) | 13 |
Total: 39 | |||
-0.1 | 204.1 | (6, 1) | 12 |
(7, 1) | 8 | ||
Total: 28 | |||
-0.1 | 287.6 | (2, 2) | 5 |
Total: 21 | |||
0.0 | 27.3 | (-14, -8) | 10 |
(11, -2) | 6 | ||
Total: 26 | |||
0.0 | 52.1 | (-9, -5) | 23 |
Total: 28 | |||
0.0 | 195.4 | (-14, -6) | 21 |
(-3, 2) | 6 | ||
Total: 34 | |||
0.0 | 247.1 | (1, 0) | 3 |
Total: 20 |
f [kHz] | () | Counts | |
0.1 | 27.6 | (-14, -8) | 6 |
(-1, -5) | 5 | ||
Total: 25 | |||
0.1 | 49.4 | (-9, -5) | 16 |
Total: 18 | |||
0.1 | 251.8 | (11, 6) | 12 |
(-15, -3) | 7 | ||
Total: 32 | |||
0.1 | 188 | (1, 0) | 7 |
Total: 23 | |||
0.2 | 28.4 | (-9, -5) | 5 |
(-1, -5) | 4 | ||
Total: 21 | |||
0.2 | 55 | (-9, -5) | 11 |
Total: 17 | |||
0.2 | 233.6 | (5, 2) | 19 |
(-7, -1) | 17 | ||
Total: 36 | |||
0.2 | 181.7 | (14, 3) | 7 |
(1, 0) | 6 | ||
(2, 0) | 3 | ||
Total: 20 |